OpenStudy (anonymous):

What is 2 and 5/6 divided by 1 and 2/9?

OpenStudy (anonymous):

ummm

OpenStudy (anonymous):

2/1 = 2 ....2/(2/9)=9 (5/6)/1=5/6=0.833 ....(5/6)/(2/9)=3.75 calculators help with this sort of stuff. just sayin.

OpenStudy (anonymous):

lol :p

OpenStudy (anonymous):

17/6 / 11/9 17/6 * 9/11 51/22

OpenStudy (anonymous):

2 and 7/22 is the final answer

OpenStudy (anonymous):

2 and 7/22 is the final answer

OpenStudy (anonymous):

HIIIIII

OpenStudy (anonymous):

eliza's answer is best although brackets would have helped. 2+(5/6)=17/6 and 1+(2/9)=11/9 So correctly it should be (17/6) / (11/9) then follow Eliza's answer